OpenStudy (anonymous):

show that \[\huge e^{sinx}-e^{-sinx}-4=0\] has no real solution

OpenStudy (rsadhvika):

say \(\large t = e^{\sin x}\)

OpenStudy (sidsiddhartha):

take \[\large e^{sinx}=y\]

OpenStudy (rsadhvika):

you will get a quadratic

OpenStudy (sidsiddhartha):

so u'll get a equation like this \[y-(1/y)-4=0\rightarrow y^2-4y-1=0\]

OpenStudy (sidsiddhartha):

now solve it

OpenStudy (anonymous):

im getting y=2+sqrt(5) and y=2-sqrt(5)

OpenStudy (sidsiddhartha):

yes take y=2+sqrt(5)

OpenStudy (sidsiddhartha):

so \[\large e^{sinx}=2+\sqrt{5}\]

OpenStudy (sidsiddhartha):

\[sinx=\log_{e} (2+\sqrt{5})\]

OpenStudy (sidsiddhartha):

again \[2+\sqrt{5}>e \rightarrow \log_{e} (2+\sqrt{5})>\log_{e}e>1\]

OpenStudy (sidsiddhartha):

hence sinx>1 which is impossible so no real solution:)

OpenStudy (anonymous):

ahh thank u :D